if 500,000 kj of heat are released by the reaction, how grams of C12H26 must have been consumed




1
Expert's answer
2022-06-03T13:32:03-0400

Mr(C12H26) = 170 g/mol

2C12H26 +37O2 = 24CO2 + 26H2O +15026 kJ


340 g C12H26 — 15026 kJ

X — 500000 kJ

X = 340 * 500000 / 15026 = 11313,7 g


ANSWER: 11313.7 g C12H26